# Aerodynamics and Aeroacoustics of da Vinci’s Aerial Screw

### v1

@@ Abstract
Leonardo da Vinci’s aerial screw, conceived in the 15th century, represents one of the earliest conceptualizations of lift-generating rotary flight. Despite its historical significance, the aerodynamic and aeroacoustic performance of this rotor has received limited scientific attention. In this study, we employ direct numerical simulations to analyze the aerodynamic forces and acoustic emissions of a modernized da Vinci aerial screw design acrossa range of Reynolds numbers (2000, 4000, and 8000). These results are compared against those from a canonical two-bladed rotor producing similar lift. The aerial screw demonstrates significantly lower mechanical power consumption and acoustic intensity per unit lift, primarily due to its larger wetted area and correspondingly lower rotational speed. Although the aerial screw exhibits a lower lift coefficient and much of its surface contributes minimally to lift generation, the net performance under iso-lift conditions highlights its efficiency and reduced noise signature. The continuous spiral geometry of the aerial screw also helps suppress blade–vortex interaction noise common in multi-bladed systems. These findings support previous scaling analyses and point toward unconventional rotor designs as viable options for low-noise aerial platforms.
@@ I Introduction
Beyond the aerodynamic performance of rotors/propellers, the rapid expansion of drones for a wide range of applications[8] has also brought to fore the importance of the aeroacoustic noise generated by rotors. There has been a concerted effort in recent times to examine the source of aeroacoustic noise from these rotors and to create new designs with a lower acoustic signatures[9]. In that regard, inspired by insect wings[10, 11] and some historical work on propeller noise reduction[12], we recently explored the simple strategy of increasing the wetted area of rotor blades for reducing aeroacoustic noise [13]. The larger wetted area allows the rotor to generate a given amount of lift at a lower rotational speed (RPM) and since aeroacoustic noise intensity scales with the second power of the rotation rate, this has been found to reduce the overall intensity of aeroacoustic noise generated for a fixed amount of lift. Interestingly, the reduction in aeroacoustic noise intensity is accompanied with a reduction in the aerodynamic power requirement, thereby providing an additional benefit for these large area rotors.
@@ II.3 Aeroacoustics Solver
[…]
@@ III.1 Cases
As described in section II.1, the rotor is modeled with a base radius of $R_{b}$, top radius of $0.5R_{b}$, a pitch of $1.31R_{b}$ and contains 1 turn. This paper will compare three cases with Reynolds number of 2000, 4000 and 8000, where the Reynolds number is described based on the tip velocity ($v_{t}$) at the base of the rotor and the base radius ($R_{b}$). The angular velocity of rotation is set as $v_{t}/R_{b}$ for the case with Reynolds number of 8000, and the angular velocity halved and quartered for the cases with Reynolds number 4000 and 2000 respectively. Note that much higher Reynolds number (such as 200,000 which would be typical for actual small size drones) are quite difficult to simulate with direct numerical simulations and we have therefore limited ourselves to comparing the designs at these lower Reynolds numbers where
Fig 2 shows the topology of the grid used in these simulations. The region around the rotor is placed in a uniform fine mesh with a gradually expanding mesh used away from the rotor body. In the plane of the rotation, a $2.4R_{b}\times 2.4R_{b}$ region centered around the rotor is finely resolved while in the vertical axis, a length of $2R_{b}$ surrounding the screw is finely resolved. Based on the grid convergence study described in the appendix A, the Reynolds number 8000 case employs a fine (32 million point) grid while those with Reynolds number of 4000 and 2000 employ a medium (13.8 million point) grid.
@@
The aim of this study is to also compare the effectiveness of aerial screw with canonical drone rotors, specifically when used to carry similar payload. We take a canonical two blade flat plate rotor from our previous work [13], with the aspect ratio of the blades equal to 5, the angle of attack is set to 20 and the distance from rotation center to outer radius is denoted using $R_{b}$. Similar to the aerial screw, the rotor is modeled as a zero thickness membrane and is simulated in a rotating reference frame. We simulate three cases with angular velocities of $v_{t}/R_{b}$, $1.51v_{t}/R_{b}$, and $2v_{t}/R_{b}$ and these cases correspond to a tip velocity based Reynolds number of 8000, 12118 and 16000 respectively. These simulations were run for a total of four rotation cycles with last two cycles used to compute the mean lift and power. As will become evident in the next section, the reason for using a higher Reynolds numbers for the two bladed rotor is to obtain a case that produces similar absolute lift as that of aerial screw for a one-to-one comparison of the performance.
@@ III.2 Flow Features
The three aerial screw cases with Reynolds number of 2000, 4000 and 8000 are simulated for a total of 6 cycles to make sure the cycle to cycle variation in the mean lift force has converged. The flow field is shown in figure 3 with the vortices shown using isosurface of Q. The case with Reynold number 2000 shows an attached leading edge vortex over almost the entire spiral leading edge. This vortex extends only short distance into the the wake before dissipating. There is also a small tip vortex that extends into the near wake. The case with Reynolds number of 4000 also shows an attached LEV that separates noticeably before the end of the spiral leading edge. A tip vortex is formed at the end of the spiral leading edge and the two vortices spiral down in the downwash of the aerial screw for later times. This case also shows the presence of a hub vortex that spirals down the hub of the aerial screw. Finally for a Reynolds number of 8000, the LEV detaches from the leading-edge a bit earlier and similar to the Re=4000 case, and we see both the detached LEV and the tip vortices spiral into the downwash. However, at this higher Reynolds number, the LEV experience an instability wherein smaller vortices are formed around the core of the large LEV. These instabilities amplify in the wake where there is a complete breakdown of the spiral vortex structures. This case also exhibits a hub vortex.
(a) Re=2000
(b) Re=4000
(c) Re=8000
Figure 3: The vortex structures shown using iso-surface of $Q$ for all the three cases and colored using the vertical velocity component. These are at a late stage in the simulation corresponding to $t/T=5.9$
@@ III.3 Aerodynamic Forces
While the lift coefficient is important in that it informs us about effectiveness of a rotor to generate lift, for practical considerations we are interested in comparing the power and aeroacoustic emission from rotors that are generating equivalent lift force (and therefore can lift the same payload). The rotors are dimensionalized using the same base radius ($R_{b}$). The lift forces for all the aerial screw and canonical rotor cases are shown in figure 6(c) and 6(d) respectively and the force values are normalized using the mean lift for the aerial screw with Reynolds number 8000 case. As one would expect, the higher Reynolds number cases with larger rotation speed generates higher lift with larger force fluctuations due to increased vortex shedding at higher Reynolds number. Among the six cases simulated here the aerial screw at Re=8000 generated roughly the same amount of mean lift as the canonical rotor at Re= 12118 (the difference is about 4%). Indeed, this match is not a coincidence since we first simulated the Re=8000 and 16000 cases for the canonical rotor and predicted the Reynolds number at which we would expect a lift equal to that of the aerial screw at Re=8000.
@@ III.4 Aeroacoustics
Figure 8: The aerial screw at Reynolds number of 8000 and the canonical rotor at Reynolds number of 12118 is compared showing (a) the directivity pattern and (b) the noise spectrum vertically above the rotation center. (c) Characteristic curve showing the acoustic intensity vs mean lift for all the aerial screw and the canonical rotor cases. The mean lift and acoustic intensity of the aerial screw at Reynolds number of 8000 is used to normalize all the values. The acoustic noise is recorded at a distance of $10R_{b}$ form the rotation center for all cases.
@@
Aerial Screw 2000 0.05 0.04 0.49 8.15×$\times 10^{-5}$
Aerial Screw 4000 0.06 0.18 0.65 8.3×$\times 10^{-3}$
@@ IV Conclusions
This study investigates the aerodynamic and aeroacoustic performance of da Vinci’s aerial screw rotor using high-fidelity direct numerical simulations. The rotor geometry was constructed based on prior design studies, and simulations were conducted across a range of Reynolds numbers to evaluate lift, mechanical power, and acoustic emissions. Our analysis confirms that while the aerial screw has a lower lift coefficient compared to a canonical two-bladed rotor, it requires significantly less mechanical power and generates substantially lower acoustic intensity for the same net lift. These advantages are primarily attributed to the screw’s large wetted area and lower rotational speed, consistent with established scaling relationships. Moreover, the continuous single-blade design of the aerial screw mitigates blade–vortex interaction noise, a key contributor to rotor aeroacoustics. While the design is not optimized for performance, the results underscore the potential benefits of unconventional rotor geometries for noise-sensitive applications. Future work could explore geometrical variations (e.g., increasing the number of turns) and extend the analysis to higher Reynolds numbers using turbulence modeling approaches. Additionally, practical considerations such as structural integrity and dynamic stability need to be examined before translating such designs into viable rotorcraft concepts.

### v2

@@ Abstract
Leonardo da Vinci’s aerial screw, conceived in the 15th century, represents one of the earliest conceptualizations of lift-generating rotary flight. Despite its historical significance, the aerodynamic and aeroacoustic performance of this rotor has received limited scientific attention. In this study, we employ direct numerical simulations to analyze the aerodynamic forces and acoustic emissions of a modernized da Vinci aerial screw design across a range of Reynolds numbers (2000, 4000, 8000, and 16000). These results are compared against those from a canonical two-bladed rotor producing similar lift. The aerial screw demonstrates 42.2% lower mechanical power consumption and 72.3% lower acoustic intensity per unit lift, primarily due to its larger wetted area and correspondingly lower rotational speed. Although the aerial screw exhibits a lower lift coefficient and much of its surface contributes minimally to lift generation, the net performance under iso-lift conditions highlights its efficiency and reduced noise signature. The continuous spiral geometry of the aerial screw also helps suppress blade–vortex interaction noise common in multi-bladed systems. These findings support previous scaling analyses and point toward unconventional rotor designs as viable options for low-noise aerial platforms.
@@ I Introduction
Beyond the aerodynamic performance of rotors/propellers, the rapid expansion of drones for a wide range of applications[8] has also brought to fore the importance of the aeroacoustic noise generated by rotors. There has been a concerted effort in recent times to examine the source of aeroacoustic noise from these rotors and to create new designs with a lower acoustic signatures. Yu [9] has listed several design modifications incorporated by traditional propellers to reduce the aeroacoustic noise. Some of these include modifying the propeller tip as in Ogee tip to prevent formation of a concentrated vortex, vane tips that have a notch at the trailing edge to produce tip vortex that are spaced apart and thus prevent blade vortex interactions, and using porous leading edge to reduce blade vortex interaction induced surface pressure fluctuation at the leading edge. Apart from conventional propeller designs, Myers [10] filled a patent for an improved “screw propeller” design in the 1890s for use in steamships. This propeller featured blades that formed a continuous loop near the tip. There are several propeller patents since then that have used looped end blades and more recently, propellers with looped end blades have reported to produce better fuel efficiency and reduced noise for marine applications [11] and better aeroacoustic performance in drones [12].
We recently explored the simple strategy of increasing the wetted area of rotor blades for reducing aeroacoustic noise [13]that was inspired by insect wings[14, 15] and some historical work on propeller noise reduction[16]. The larger wetted area allows the rotor to generate a given amount of lift at a lower rotational speed (RPM) and since aeroacoustic noise intensity scales with the second power of the rotation rate, this has been found to reduce the overall intensity of aeroacoustic noise generated for a fixed amount of lift. Interestingly, the reduction in aeroacoustic noise intensity is accompanied with a reduction in the aerodynamic power requirement, thereby providing an additional benefit for these large area rotors.
@@ II.3 Aeroacoustics Solver
$\begin{split}4\pi p^{\prime}({\bf x},t)=&\frac{1}{c}\int\Big{[}\frac{\dot{L}_{r}}{r(1-M_{r})^{2}}\Big{]}_{t-r/c}dS\\ &+\int\Big{[}\frac{L_{r}-L_{M}}{r^{2}(1-M_{r})^{2}}\Big{]}_{t-r/c}dS\\ &+\frac{1}{c}\int\Big{[}\frac{L_{r}(r\dot{M_{r}}+c(M_{r}-M^{2}))}{r^{2}(1-M_{r})^{3}}\Big{]}_{t-r/c}dS\,\,.\end{split}$ (6)
@@ III.1 Cases
As described in section II.1, the rotor is modeled with a base radius of $R_{b}$, top radius of $0.5R_{b}$, a pitch of $1.31R_{b}$ and contains 1 turn. This paper will compare four cases with Reynolds number of 2000, 4000, 8000, and 16000, where the Reynolds number is described based on the tip velocity ($v_{t}$) at the base of the rotor and the base radius ($R_{b}$). The angular velocity of rotation is set as $v_{t}/R_{b}$ for the case with Reynolds number of 8000, and the angular velocity halved, quartered, and doubled for the cases with Reynolds number 4000, 2000, and 16000 respectively. Note that much higher Reynolds number (such as 200,000 which would be typical for actual small size drones) are quite difficult to simulate with direct numerical simulations and we have therefore limited ourselves to comparing the designs at these lower Reynolds numbers.
Fig 2 shows the topology of the grid used in these simulations. The region around the rotor is placed in a uniform fine mesh with a gradually expanding mesh used away from the rotor body. In the plane of the rotation, a $2.4R_{b}\times 2.4R_{b}$ region centered around the rotor is finely resolved while in the vertical axis, a length of $2R_{b}$ surrounding the screw is finely resolved. Based on the grid convergence study described in the appendix A, the Reynolds number 8000 and 16000 case employs a fine (32 million point) grid while those with Reynolds number of 4000, and 2000 employ a medium (13.8 million point) grid.
@@
The aim of this study is to also compare the effectiveness of aerial screw with canonical drone rotors, specifically when used to carry similar payload. We take a canonical two blade flat plate rotor from our previous work [13], with the aspect ratio of the blades equal to 5, the angle of attack is set to 20 and the distance from rotation center to outer radius is denoted using $R_{b}$. Similar to the aerial screw, the rotor is modeled as a zero thickness membrane and is simulated in a rotating reference frame. We simulate three cases with angular velocities of $v_{t}/R_{b}$, $1.51v_{t}/R_{b}$, and $2v_{t}/R_{b}$ and these cases correspond to a tip velocity based Reynolds number of 8000, 12118 and 16000 respectively. The input parameters corresponding to all 7 cases are summarized in table 1. These simulations were run for a total of four rotation cycles with last two cycles used to compute the mean lift and power. As will become evident in the next section, the reason for using a higher Reynolds numbers for the two bladed rotor is to obtain a case that produces similar absolute lift as that of aerial screw for a one-to-one comparison of the performance.
Table 1: The input Reynolds number and the rotation speed for all the seven cases.
Rotor Re $\Omega$
Aerial Screw 2000 $0.25v_{t}/R_{b}$
Aerial Screw 4000 $0.5v_{t}/R_{b}$
Aerial Screw 8000 $v_{t}/R_{b}$
Aerial Screw 16000 $2v_{t}/R_{b}$
Canonical Rotor 8000 $v_{t}/R_{b}$
Canonical Rotor 12118 $1.51v_{t}/R_{b}$
Canonical Rotor 16000 $2v_{t}/R_{b}$
@@ III.2 Flow Features
The four aerial screw cases with Reynolds number of 2000, 4000, 8000, and 16000 are simulated for a total of 6 cycles to make sure the cycle to cycle variation in the mean lift force has converged. The flow field is shown in figure 3 with the vortices shown using isosurface of Q. The case with Reynold number 2000 shows an attached leading edge vortex over almost the entire spiral leading edge. This vortex extends only short distance into the the wake before dissipating. There is also a small tip vortex that extends into the near wake. The case with Reynolds number of 4000 also shows an attached LEV that separates noticeably before the end of the spiral leading edge. A tip vortex is formed at the end of the spiral leading edge and the two vortices spiral down in the downwash of the aerial screw for later times. This case also shows the presence of a hub vortex that spirals down the hub of the aerial screw. Finally for a Reynolds number of 8000, the LEV detaches from the leading-edge a bit earlier and similar to the Re=4000 case, and we see both the detached LEV and the tip vortices spiral into the downwash. However, at higher Reynolds numbers of 8000 and 16,000, the LEV experience an instability wherein smaller vortices are formed around the core of the large LEV. These instabilities amplify in the wake where there is a complete breakdown of the spiral vortex structures. This case also exhibits a hub vortex.
Figure 3: The vortex structures shown using iso-surface of $Q$ for Reynolds number (a) 2000, (b) 4000, (c) 8000, and (d) 16000 cases and colored using the vertical velocity component. These are at a late stage in the simulation corresponding to $t/T=5.9$
@@ III.3 Aerodynamic Forces
While the lift coefficient is important in that it informs us about effectiveness of a rotor to generate lift, for practical considerations we are interested in comparing the power and aeroacoustic emission from rotors that are generating equivalent lift force (and therefore can lift the same payload). The rotors are dimensionalized using the same base radius ($R_{b}$). The lift forces for all the aerial screw and canonical rotor cases are shown in figure 6(c) and 6(d) respectively and the force values are normalized using the mean lift for the aerial screw with Reynolds number 8000 case. As one would expect, the higher Reynolds number cases with larger rotation speed generates higher lift with larger force fluctuations due to increased vortex shedding at higher Reynolds number. Among the seven cases simulated here the aerial screw at Re=8000 generated roughly the same amount of mean lift as the canonical rotor at Re= 12118 (the difference is about 4%). Indeed, this match is not a coincidence since we first simulated the Re=8000 and 16000 cases for the canonical rotor and predicted the Reynolds number at which we would expect a lift equal to that of the aerial screw at Re=8000.
@@ III.4 Aeroacoustics
Figure 8: The aerial screw at Reynolds number of 8000 and the canonical rotor at Reynolds number of 12118 is compared showing (a) the directivity pattern and (b) the noise spectrum vertically above the rotation center. (c) Characteristic curve showing the acoustic intensity vs mean lift for all the aerial screw and the canonical rotor cases. The mean lift, mean directivity, and acoustic intensity of the aerial screw at Reynolds number of 8000 is used to normalize all the values. The acoustic noise is recorded at a distance of $10R_{b}$ form the rotation center for all cases.
@@
[…]
@@
III.5 Noise Source Analysis
In the previous section III.4, we discussed the aeroacoustic noise generated at a distance of $10R_{b}$ from the rotation center and we note in figure 8a that the aeroacoustic noise is the largest in the vertical direction. Since the noise at a far-field location is also important in the context of drone rotors [8] and the drone rotor noise in the vertical direction are larger than noise in rotation plane and will travel directly to the ground to create annoyance, in this section we will identify and analyze the vortical structures that are primarily responsible for the far-field noise in vertical (or lift) direction.
At low Mach numbers, the aerodynamic force on the rotor surface, $\mathbf{L}$ is the major source of noise[31] (see Eq.6) and we use the acoustic partitioning method that was proposed by by Seo, Menon, and Mittal [32] who used the Ffowcs Williams - Hawkings equation [31] written for a compact source to calculate the far-field noise generated by the forces acting on the body ($\mathbf{F}_{R}=\int{\mathbf{L}}dS$) as,
$p^{{}^{\prime}}=\frac{1}{4\pi}\left(\frac{{\bf\dot{F_{R}}\cdot\hat{r}}}{cr}+\frac{{\bf F_{R}\cdot\hat{r}}}{r^{2}}\right)_{t-\frac{r}{c}}\,\,$, (15)
where $r$ is the distance to the recording location, $^{\bf\hat{r}}$ is the unit normal pointing from source to the recording location and $c$ is the speed of the sound. Note that when the recording location is in the lift direction, ${\bf F_{R}\cdot\hat{r}}=F_{L}$, where $F_{L}$ is the lift force generated by the rotor. In the acoustic partitioning method, the force is further decomposed by using the force partitioning method [33, 34, 35] (FPM) into four components, namely the force due to vortices, added mass effect, viscous diffusion of the momentum, and due to the acceleration of the body. This is done by first solving the Laplace equation for the influence field ($\phi$) given by,
[…]
where, $B$ and $\Sigma$ are the body and domain boundary respectively, $V$ is the domain volume, and $n_{z}$ is the normal in the lift direction since we are decomposing the lift force. The gradient of this influence field is then projected onto the Navier-Stokes equation and the derivation along with the details of the FPM can be found in Menon and Mittal [33, 34, 35]. For the present rotor flows, we have found that the major force contribution comes from the force due to the vortices. For instance, if we consider the aerial screw, the vortex induced lift force accounts for 96% of the net lift force while the force due to the viscous diffusion of the momentum accounts for only 4% of the net lift force and the remaining two force terms are 0. The reasoning for this can be found in Prakhar, Seo, and Mittal [36]. Thus, for the present study, the vortex induced lift force is a major source of aeroacoustic sound generation. This vortex induced lift force ($F_{Q}$) is defined as,
$F_{Q}=\int_{V}-2\rho\phi QdV\,\,=\int_{V}f_{Q}dV,$ (17)
[…]
$p_{Q}^{{}^{\prime}}=\frac{1}{4\pi}\left(\frac{{\bf\dot{F_{Q}}\cdot\hat{r}}}{cr}+\frac{{\bf F_{Q}\cdot\hat{r}}}{r^{2}}\right)_{t-\frac{r}{c}}\,\,$, (18)
Some applications of this method can be found in Seo, Menon, and Mittal [32], Seo et al. [38], Prakhar, Seo, and Mittal [39].
Here, we employ this method to represent the aeroacoustic noise at a given far-field location into a volume source contribution where all the vortical structures are associated with the portion of the noise they generate. For this, we use the equation 17 and rewrite the vortex induced sound pressure (equation 18) as,
$p_{Q}^{{}^{\prime}}=\int_{V}\left[\frac{1}{4\pi}\left(\frac{{\bf\dot{f_{Q}}\cdot\hat{r}}}{cr}+\frac{{\bf f_{Q}\cdot\hat{r}}}{r^{2}}\right)_{t-\frac{r}{c}}\right]dV=\int_{V}\hat{p_{Q}}dV\,\,$, (19)
where $\hat{p_{Q}}$ describes the “specific sound pressure” at each point in the domain volume, and integrating this over the entire domain will yield the total sound pressure ($p_{Q}^{{}^{\prime}}$). Thus, spatial distribution of $\hat{p_{Q}}$ indicates the major source region or the vortical structure most responsible for the noise generation. The volume integral can also be restricted to a specific vortical structure to get the far-field noise due to a specific vortex as well. We will use this specific sound pressure to color the vortices and identify the vortical structures that are responsible for most of the noise generation.
We apply this method to both the aerial screw and the canonical rotor at Reynolds number of 8000 and 12118 respectively (i.e. when both these rotors produce similar lift force). The recording location is set at a distance of $50R_{b}$ in the vertical direction measured from the rotation center since we need larger recording distance to keep the point source assumption valid.
Figure 9: The flow field is shown using Q iso-surface and colored using (a) $f_{Q}$, and (b) $\hat{p_{Q}}$ for the aerial screw at t/Tt/T=5.89 . Similarly flow field for the canonical rotor colored using (c) $f_{Q}$, and (d) $\hat{p_{Q}}$ is shown at t/Tt/T=4.0. The integral of $\hat{p_{Q}}$ will give the total sound pressure recorded at at time $(t+r/c)/T$ on the recording location, where $r/c$ accounts for time delay required for propagation of sound waves.
The vortex induced lift force density is shown for the aerial screw and the canonical rotor in figures 9a and 9c respectively and we note that the entire spiral outer edge of the screw produces a major portion of the lift force while for the canonical rotor, this region is located along the leading edge and near the tip. The specific sound pressure for the aerial screw (figure 9b) reveals that while the spiral leading edge vortex generates greater amount of lift compared to smaller shed “c” shaped vortices near the base of the screw, this leading edge vortex provide temporally invariant lift and hence do not contribute much to the acoustic noise. The shed vortices in the wake and downwash region, while they are not large contributors to the lift force, they create larger pressure fluctuations on the rotor surface as these vortices shed and thus they are responsible for most of the acoustic noise. We contrast this to the specific sound pressure of the canonical rotor shown in figure 9d and we note that the leading edge vortex that is responsible for most of the lift force and since they also temporally vary in magnitude (this was also verified using the time derivative of the flow field), the leading edge vortex is also a significant source of the noise. The shedding at the tip and a small region in the wake of the tip also contributes to the acoustic noise since the shedding would create pressure fluctuations on the rotor surface. The shed vortical structures in the wake away from the rotor surface does not produce much lift force or the associated aeroacoustic noise since the influence field ($\phi$) decays rapidly away from the surface.
@@ IV Conclusions
This study investigates the aerodynamic and aeroacoustic performance of da Vinci’s aerial screw rotor using high-fidelity direct numerical simulations. The rotor geometry was constructed based on prior design studies, and simulations were conducted across a range of Reynolds numbers to evaluate lift, mechanical power, and acoustic emissions. Our analysis confirms that while the aerial screw has a lower lift coefficient compared to a canonical two-bladed rotor, it requires significantly less mechanical power and generates substantially lower acoustic intensity for the same net lift. These advantages are primarily attributed to the screw’s large wetted area and lower rotational speed, consistent with established scaling relationships. Moreover, the continuous single-blade design of the aerial screw mitigates blade–vortex interaction noise, a key contributor to rotor aeroacoustics. We also note that the temporally invariant magnitude of the leading edge vortex of the screw helps to generate lift without significant contribution to the far-field noise. While the design is not optimized for performance, the results underscore the potential benefits of unconventional rotor geometries for noise-sensitive applications. Future work could explore geometrical variations (e.g., increasing the number of turns) and extend the analysis to higher Reynolds numbers using turbulence modeling approaches. We expect similar improvement in mechanical power and acoustic intensity for the higher Reynolds number aerial screw simulations as well since the lift coefficient has already converged for both the aerial screw and the canonical rotor cases presented here and therefore the scaling law presented in section II.4 should hold. Additionally, practical considerations such as structural integrity and dynamic stability need to be examined before translating such designs into viable rotorcraft concepts.